You are assigned as a stunt coordinator for action movie. They are trying to do a stunt where a car leaps over a lake that is 200m wide from the top of a cliff that is 100m high. There is no ramp, the car will go off a flat cliff. The maximum speed of the car is 30m/s. Will the car clear the lake? If not, how many meters of road needs to be built over the lake to catch the car? Please show which formula to use and all your work.
Answer #1

Time of flight can be found using second equation of motion in vertical,

h = 0.5gt^2

t = sqrt(2h/g) = sqrt(2*100/9.8) = 4.5175 s

Maximum horizontal distance = v*t = 30*4.5175 = 135.5 m

maximum horizontal distance of car is less than width of lake, so car will not clear the lake.

Minimum distance roads to be built = 200m - 135.5 m = 64.5 m answer